OpenStudy (anonymous):

help me to factor this: x^2-27x^2a^3

OpenStudy (anonymous):

You can factorize by x^2: x^2*(1 - 27a^3) The difference within brackets is a difference of cubes: a^3 - b^3 = (a-b)(a^2 + ab + b^2) Let a = 1 and b = 3a 1 - 27a^3 = (1-3a)(1 + 3a + 9a^2) The factors are: x^2*(1 - 3a)(1 + 3a + 9a^2)

OpenStudy (anonymous):

First you want to factor out anything you see in common between the two terms. Looks like there is an x^2 in common: \[x^2-27x^2a^3 \Rightarrow x^2(1-27a^3)\] Then, you should notice something about that last term: \[1-27a^3 \Rightarrow 1^3-(3a)^3\] That is a difference of 2 cubes, which follows this formula: \[x^3-y^3 = (x-y)(x^2+xy+y^2)\] so we have: \[1^3-(3a)^3 = (1-3a)(1+3a+9a^2)\] So the final answer is: \[x^2(1-3a)(1+3a+9a^2)\]